OpenStudy (anonymous):

simplify and reduce the expression to lowest terms. 22x/7x+2-15x-2/7x+2

OpenStudy (anonymous):

better use lots of brackets..............

OpenStudy (anonymous):

this is jiberish to me

OpenStudy (anonymous):

do you want me to rewrite it

OpenStudy (saifoo.khan):

The answer is 1.

OpenStudy (anonymous):

can you show the work?

OpenStudy (saifoo.khan):

\[\frac{22x-15x-2}{7x+2}= \frac{7x+2}{7x+2}=1\]

OpenStudy (anonymous):

Thanks

OpenStudy (saifoo.khan):

uw.